A cube of volume 1 cubic metres and specific gravity 2 is very slowly being lowered in to a lake with help of a massless string tied to one of its vertices A (i.e., body diagonal AB is perpendicular to surface of water).

Initially, the other opposite vertex B is just touching the surface of water and finally, vertex A is just beneath the surface of water. (Ignore change in KE of water). (g= 10ms2).

Find
(a) Work done by gravity on cube.
(b) Work done by gravity on water
(c) Work done by gravity on water + cube system.
(d) Work done by tension on cube.
1
Expert's answer
2020-08-25T11:13:42-0400

a) By the formula of work:

"A=P*S"

where P is ponderosity of cube.

Specific gravity: "\gamma=\frac{P}{V}->P=\gamma*V"

"A=\gamma*V*S"

"S=d", "d" is the diagonal of cube. If "a=1 ->d=a\sqrt{\smash[b]{3}}=\sqrt{\smash[b]{3}}\approx1.7"

"A=2*1*1.7=3.4" (J)


b) Work done by gravity of water:

"A=F_A*S"

"F_A" is extrude Archimedes force. "F_A=\rho*g*V=1000*10*1=10000" (N)

"A=10000*1.7=17000" (J)


c) Work done by gravity on water + cube system:

"A=17000+3.4=17003.4" (J)


d) Work done by tension on cube:

"A=3.4" (J)